# 22affiliationtext: Booth School of Business

### v1

@@ Acknowledgements
This research was supported in part through the computational resources and staff contributions provided for the Mercury high performance computing cluster at The University of Chicago Booth School of Business which is supported by the Office of the Dean.
Appendix A Identifiability of the proposed model
@@ A.1 Set-up
The random variables $\mathbf{c}_{j}$ will be referred to as “concepts”: Intuitively they represent abstract concepts such as shape, size, colour, etc. that may not be perfectly represented by any particular embedding in $\mathbf{e}$. The linear representation hypothesis (Section 3) is an empirical observation that abstract concepts can be approximately represented as linear projections of a sufficiently flexible latent space. This is operationalized by assuming that $\mathbf{c}_{j}\approx C_{j}\mathbf{e}$ (see (A1) below). The matrix $C_{j}$ is referred to as the representation of the concept $\mathbf{c}_{j}$. If $\mathbf{c}_{j}\in\mathbb{R}$ (i.e. $C_{j}\in\mathbb{R}^{1\times d_{e}}$), then $\mathbf{c}_{j}$ is referred to as an atom and $C_{j}$ its atomic representation. We will often denote atoms by $\mathbf{a}_{j}$ for notational clarity.
@@
Note that these assumptions are not necessary in practice—as evidenced by our experiments in Section 5 which clearly violate these assumptions—and are merely used here to provide a proof-of-concept identifiability result based on our architectural choices.
Our goal is to identify the concept marginal $p(\mathbf{c})$ and concept representations $(C_{1},\ldots,C_{d_{c}})$ from concept interventions. Concept interventions are well-defined through the structural causal model defined by the DAG $\mathsf{G}$ and its structural equations. We consider single-node concept interventions where $\mathbf{c}_{j}$ is stochastically set to be centered at $\boldsymbol{\mu}_{j}$ with variance $\Omega_{j}\succ 0$ (making this precise requires some set-up; see Appendix A.3.1 for details).
@@
This definition of identifiability, which aligns with similar notions of identifiability that have appeared previously (e.g. Squires et al., 2023; von Kügelgen et al., 2023; Rajendran et al., 2024; Fokkema et al., 2025), implies in particular that the concepts $\mathbf{c}_{j}$ are identifiable up to permutation, shift, and scale, and that their representations are identifiable up to a linear transformation. As a result, the concept marginal $p(\mathbf{c})$ is also identifiable (again up to permutation, shift, and scale as well). The shift ambiguity can of course be resolved by assuming the concepts are centered (zero-mean), and the scale ambiguity can be resolved by assuming concepts are normalized (e.g. unit norm).
@@
Assumption 1 (Atomic representations).
There exists a set of atomic representations $\mathcal{A}=\{\mathbf{a}_{1},\ldots,\mathbf{a}_{n}\}$ of linearly independent vectors $\mathbf{a}_{j}\in\mathbb{R}^{d_{e}}$ such that the rows of each representation $C_{j}$ are in $\mathcal{A}$. We denote the indices of $\mathcal{A}$ that appear as rows of $C_{j}$ by $S^{j}$, i.e. $S^{j}=\{i:\mathbf{a}_{i}\text{ is a row in $C_{j}$}\}$. We assume that $\cup_{j}S^{j}=\mathcal{A}$, i.e. every atom in $\mathcal{A}$ appears in some concept representation $C_{j}$.
@@
$B_{ij}=\begin{cases}\frac{(\boldsymbol{\mu}_{j})_{k}}{\omega_{j}^{2}}&\text{if $i\in S^{j}$ and the $k$th row of $C_{j}$ is $\mathbf{a}_{i}$,}\\ 0&\text{otherwise.}\end{cases}$ (29)
@@
Assumption 2 (Concept diversity).
@@
A.2 Formal statement of Theorem 4.1
@@
Theorem A.1.
Under Assumptions 1-2 and given single-node interventions on each concept $\mathbf{c}_{j}$, we can identify the representations $C_{j}$ and the latent concept distribution $p(\mathbf{c})$, in the sense of (26-27).
In the context of learning disentangled representations, identifying the concept marginal $p(\mathbf{c})$ is particularly relevant, since this implies we also learn certain (in)dependence relationships between the concepts. In other words, if the “true” concept representations are disentangled (i.e. independent), then we identify disentangled concepts.
@@
  • While our proof ultimately relies on techniques from Rajendran et al. (2024), Theorem A.1 does not trivially follow from their results, which do not cover interventions over the concept marginal $p(\mathbf{c})$. Extending their results to the case of concept interventions, including making sure these interventions are well-defined causal objects, is the main technical difficulty in our proof. This is surprisingly tricky since we do not assume a full structural causal model over $(\mathbf{x},\mathbf{c},\mathbf{e})$: Interventions are only well-defined over $\mathbf{c}$, which makes deducing the downstream effects on the observed $\mathbf{x}$ somewhat subtle.
  • Various results in CRL (e.g. Buchholz et al., 2023; von Kügelgen et al., 2023; Varici et al., 2024, see Section 3 for more references) prove identifiability results under interventions, however, these results require interventions directly on the embeddings $\mathbf{e}$, as opposed to the concepts $\mathbf{c}$. Theorem A.1, by contrast, requires only $d_{c}\ll d_{e}$ total interventions and in doing so directly address the technical challenges associated with intervening directly on concepts as opposed to embeddings.
@@ A.3.1 Concept interventions
Lemma A.3.
For any distribution $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, the interventional distribution $P^{j}$ defined by (31) is also an element of $\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, i.e. $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$.
@@
Proof.
Lemma A.3 implies that $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Thus Lemma A.2 implies the existence of joint distributions $\widetilde{P},\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ such that
$\widetilde{p}(\mathbf{c})=p(\mathbf{c})\quad\text{ and }\quad\widetilde{p}^{j}(\mathbf{c})=p^{j}(\mathbf{c})\text{ for all }\mathbf{c}\in\mathcal{C}.$ (32)
(1) follows from the definition of $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ for $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ from Lemma A.2 above; i.e. (32) above.
@@
$\begin{split}\widetilde{p}^{j}(\mathbf{x},\mathbf{c},\mathbf{e})&=\widetilde{p}^{j}(\mathbf{x}\mid\mathbf{e})\widetilde{p}^{j}(\mathbf{c}\mid\mathbf{e})\widetilde{p}^{j}(\mathbf{e}),\\ &=\widetilde{p}^{j}(\mathbf{x}\mid\mathbf{e}){p}^{j}(\mathbf{c}\mid\mathbf{e})\widetilde{p}^{j}(\mathbf{e}),\\ \end{split}$
@@
(3) follows similarly to (2). Namely, since the intervention only perturbs the conditional factors specified above, we can always choose $f^{j}=f$. ∎
@@
Proposition A.5.
Let $I_{j}:=\{j\}$ be the collection of single-node intervention targets, and consider the interventional distributions $P^{1},\ldots,P^{d_{c}}$ for $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Then there exists $\widetilde{P},\widetilde{P}^{1},\ldots,\widetilde{P}^{d_{c}}\in\mathcal{M}(\mathsf{G})$ such that
@@
Proof.
Fix a choice of $\widetilde{P}\in\mathcal{M}(\mathsf{G})$ for $P$, which specifies a function $f$. For the chosen $\widetilde{P}$, we construct $\widetilde{P},\widetilde{P}^{1},\ldots,\widetilde{P}^{d_{c}}$ according to Proposition A.4. The result follows by applying Proposition A.4 for every $j$. ∎
Proposition A.5 will allow us (below) to define more formally the environments that are used to identify concepts. Intuitively, each environment corresponds to single-node interventions on a single concept $\mathbf{c}_{j}$, however, we can only observe the effect this intervention has on the observed $\mathbf{x}$. The purpose of Propositions A.4-A.5 are to trace the dependence between $\mathbf{x}$ and $\mathbf{c}$—via the embeddings $\mathbf{e}$—according to the model (A1)-(A3).
Formally, let $I_{j}=\{j\}$ with $\eta_{j}^{\prime}\sim\mathcal{N}(0,\omega_{j}^{2})$ and $\mathbf{c}_{j}=\boldsymbol{\mu}_{j}+\eta_{j}^{\prime}$ according to (31), i.e. each environment corresponds to a single-node intervention on the $j$th concept $\mathbf{c}_{j}$. By Proposition A.5, there exist post-intervention joint distributions $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ over $(\mathbf{x},\mathbf{c},\mathbf{e})$ such that
@@
and $f^{j}=f$ for all $j$. Then the $j$th environment corresponds to sampling $\mathbf{x}=f(\mathbf{e})$ where $\mathbf{e}\sim p_{j}(\mathbf{e}):=\widetilde{p}^{j}(\mathbf{e}\,|\,\mathbf{c}_{j}=\boldsymbol{\mu}_{j})$.
A.3.2 Reduction to concept identification
We begin by computing the log-likelihood ratio between the observational and interventional environments. The idea is that the concept representations are revealed through fluctuations in this likelihood ratio. Using Proposition A.5, we have
@@ A.3.2 Reduction to concept identification
$h_{j}(\mathbf{e}):=\log p_{0}(\mathbf{e})-\log p_{j}(\mathbf{e})$ $\propto\sum_{i=1}^{n}\frac{1}{2}M_{ij}\mathbf{e}_{i}^{2}-B_{ij}\mathbf{e}_{i}+O(1).$ (36)
These functions are convex, which allows us to identify the $|S_{T}|$ where $S_{T}=\cup_{j\in T}S_{j}$ by minimizing the convex function $\sum_{j\in T}h_{j}(\mathbf{e})$. Then a similar induction argument identifies the matrices $M$ and $B$. Now, given two distinct representations of $p(\mathbf{x})$ via nonlinear mixing functions $f$ and $\widetilde{f}$, define $\varphi=\widetilde{f}^{-1}\circ f$. Let $\mathbf{h}=(h_{1},\ldots,h_{n})$ and write $\mathbf{e}^{c}=(\mathbf{e}_{1},\ldots,\mathbf{e}_{n})$ and $\mathbf{e}=(\mathbf{e}^{c},\mathbf{e}^{\perp})\in\mathbb{R}^{n\times(d_{c}-n)}$. Let $\iota^{\perp}(\mathbf{e}^{c})=(\mathbf{e}^{c},0)$, $\pi^{c}(\mathbf{e})=\mathbf{e}^{c}$, and $\varphi^{\perp}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ by $\varphi^{\perp}(\mathbf{e}^{c})_{i}=\varphi(\mathbf{e},0)_{i}$. Note that via the change of variables formula, we can write $\mathbf{h}(\mathbf{e})=\mathbf{H}(\mathbf{x})$ for some function $\mathbf{H}$. Then
@@
which is the crucial relation used in the proof of Theorem 1 of Rajendran et al. (2024). From here, identifiability follows from a straightforward but lengthy linear algebraic argument based on the first-order conditions for minimizing $\mathbf{h}$ and the identifiability of $M,B$ proved above. The proof of Theorem A.1 follows.
Appendix B quad: A semi-synthetic benchmark for OOD composition
@@
These properties make quad especially useful for both quantitative and qualitative evaluation; indeed, it was created precisely for this reason. Example images are shown in Figure 5.
To describe our experimental setting, let $\mathbf{c}_{j}$ indicate the $j$th concept as listed above. We used seven different contexts, including an observational context and one context each for single-concept interventions on $(\mathbf{c}_{1},\ldots,\mathbf{c}_{6})$, i.e. quad1, quad2, quad3, quad4, size, and orientation. We use a sample size of 50 000 images per context. In each context, $\mathbf{c}_{7}$ (object) and $\mathbf{c}_{8}$ (shape) are sampled uniformly at random. The remaining six concepts were sampled as follows:
@@
Similarly, we can compose multiple interventions together, as in Figure 6. This facilitates evaluation of the concept composition capabilities of models trained on single interventions, providing a ground truth distribution to compare against generated OOD samples like in Figure 7.
Figure 5: Example images obtained from the quad dataset. Each subfigure corresponds to a different context, indicated by the title (obs = observational; the others indicate a single-node intervention).
@@
For comparison, see Figure D.1.2 for examples of generated multi-concept compositions (OOD) from a simple 3-layer convolutional model. The training data did not contain any multi-concept compositions.
Figure 6: Example images obtained from the quad dataset. Each subfigure corresponds to a different double-concept intervention context, indicated by the title.
Appendix C Experimental details
In this appendix we provide additional details on our experimental protocol and set-up.
C.1 Evaluation metrics

### v2

@@ Acknowledgments
The work by AM was supported by Novo Nordisk Foundation Grant NNF20OC0062897. This research was supported in part through the computational resources and staff contributions provided for the Mercury high performance computing cluster at The University of Chicago Booth School of Business which is supported by the Office of the Dean.
Supplementary material:
Intervening to learn and compose
causally disentangled representations
Appendix A Related work
Our work is closely related to several parallel lines of work on structured generative models, causal representation learning, linear representations, and OOD generalization. Below we compare and contrast our contributions against this literature.
Structured generative models
To provide structure such as hierarchical, graphical, causal, and disentangled structures as well as other inductive biases in the latent space, there has been a trend towards building structured generative models that directly impose this structure a priori. Early work looked at incorporating fixed, known structure into generative models, such as autoregressive, graphical, and hierarchical structure (Germain et al., 2015; Johnson et al., 2016; Sønderby et al., 2016; Webb et al., 2018; Weilbach et al., 2020; Ding et al., 2021; Mouton and Kroon, 2023). This was later translated into known causal structure (Kocaoglu et al., 2017; Markham et al., 2023). When the latent structure is unknown, several techniques have been developed to learn useful (not necessarily causal) structure from data (Li et al., 2018; He et al., 2019; Wehenkel and Louppe, 2021; Kivva et al., 2022; Moran et al., 2022). More recently, based on growing interest in disentangled (Bengio, 2013) and/or causal (Schölkopf et al., 2021) representation learning, methods that learn causal structure have been developed (Markham and Grosse-Wentrup, 2020; Moraffah et al., 2020; Yang et al., 2021; Ashman et al., 2022; Shen et al., 2022; Kaltenpoth and Vreeken, 2023). In contrast to this line of work, which emphasizes hard graphical constraints, we neither learn nor impose any fixed graphical structure. Instead, we emphasize performance on concept learning and composition as downstream tasks.
Causal disentanglement and representation learning
Causal representation learning (Schölkopf et al., 2021) is a rapidly developing area that involves, among other goals, two key objectives: 1) Learning disentangled latent factors along with 2) Learning latent causal structure. The former is what we refer to as causal disentanglement (see Definition 2.1 and its causal interpretation). The latter problem has seen tremendous theoretical progress in recent years (Brehmer et al., 2022; Shen et al., 2022; Lachapelle et al., 2022; Moran et al., 2022; Kivva et al., 2022; Buchholz et al., 2023; Gresele et al., 2021; Ahuja et al., 2023; von Kügelgen et al., 2023; Morioka and Hyvärinen, 2023; Liu et al., 2023; Mameche et al., 2023; Yao et al., 2023; Varici et al., 2023; Sturma et al., 2023; Xu et al., 2024; Zhang et al., 2024; Li et al., 2024; Varici et al., 2024; Bing et al., 2024; Ahuja et al., 2024; Talon et al., 2024; Yao et al., 2024). Recent work has also pushed in the direction of identifying concepts (Leemann et al., 2023; Rajendran et al., 2024; Fokkema et al., 2025). Our work draws inspiration from these lines of work, which articulate precise conditions under which latent causal discovery is possible in principle. By contrast, our focus is on methodological aspects of causal disentanglement in practice. While the former is a notoriously difficult task, the latter only needs to exploit learned causal invariances: In particular, causal disentanglement is possible without learning latent causal structure.
Linear representations
Generative models are known to represent concepts linearly in embedding space (e.g. Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015); see also Levy and Goldberg (2014); Arora et al. (2016); Gittens et al. (2017); Allen and Hospedales (2019); Ethayarajh et al. (2018); Seonwoo et al. (2019). This phenomenon has been well-documented over the past decade in both language models (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016; Conneau et al., 2018; Tenney et al., 2019; Elhage et al., 2022; Burns et al., 2022; Tigges et al., 2023; Nanda et al., 2023; Moschella et al., 2022; Li et al., 2023; Park et al., 2023; Gurnee et al., 2023; Cunningham et al., 2024; Jiang et al., 2024) and computer vision (Radford et al., 2015; Raghu et al., 2017; Bau et al., 2017; Engel et al., 2017; Kim et al., 2018; Chen et al., 2020; Wang et al., 2023; Trager et al., 2023). Our approach actively exploits this tendency by searching for concept representations as linear projections of the embeddings learned by a black-box model. In contrast to this prior work, our emphasis is on applying these empirical observations for causal disentanglement, compositional generalization, and OOD generation.
OOD generalization
A growing line of work studies the OOD generalization capabilities of generative models, with the general observation being that existing methods struggle to generalize OOD (Xu et al., 2022; Mo et al., 2024; Montero et al., 2022, 2021; Schott et al., 2022). It is worth noting that most if not all of this work evaluates OOD generalization using reconstruction on held-out OOD samples, as opposed to generation. For example, a traditional VAE may be able to reconstruct held-out samples, but it is not possible to actively sample OOD. See Section 2 for more discussion.
Appendix B Identifiability of the proposed model
@@ B.1 Set-up
The random variables $\mathbf{c}_{j}$ will be referred to as “concepts”: Intuitively they represent abstract concepts such as shape, size, colour, etc. that may not be perfectly represented by any particular embedding in $\mathbf{e}$. The linear representation hypothesis (Section A) is an empirical observation that abstract concepts can be approximately represented as linear projections of a sufficiently flexible latent space. This is operationalized by assuming that $\mathbf{c}_{j}\approx C_{j}\mathbf{e}$ (see (A1) below). The matrix $C_{j}$ is referred to as the representation of the concept $\mathbf{c}_{j}$. If $\mathbf{c}_{j}\in\mathbb{R}$ (i.e., $C_{j}\in\mathbb{R}^{1\times d_{e}}$), then $\mathbf{c}_{j}$ is referred to as an atom and $C_{j}$ its atomic representation. We will often denote atoms by $\mathbf{a}_{j}$ for notational clarity.
@@
Note that these assumptions are not necessary in practice—as evidenced by our experiments in Section 4 which clearly violate these assumptions—and are merely used here to provide a proof-of-concept identifiability result based on our architectural choices. The representation (5) implies, in particular, the usual factorization
$p(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})=\prod_{j=1}^{d_{c}}p(\mathbf{c}_{j}\,|\,\operatorname{pa}_{\mathsf{G}}(j)),$
which formalizes the notion of causal disentanglement as in Schölkopf et al. (2021).
Our goal is to identify the concept marginal $p(\mathbf{c})$ and concept representations $(C_{1},\ldots,C_{d_{c}})$ from concept interventions. Concept interventions are well-defined through the structural causal model defined by the DAG $\mathsf{G}$ and its structural equations. We consider single-node concept interventions where $\mathbf{c}_{j}$ is stochastically set to be centered at $\boldsymbol{\mu}_{j}$ with variance $\Omega_{j}\succ 0$ (making this precise requires some set-up; see Appendix B.3.1 for details).
@@
This definition of identifiability, which aligns with similar notions of identifiability that have appeared previously (e.g. Squires et al., 2023; von Kügelgen et al., 2023; Rajendran et al., 2024; Fokkema et al., 2025), implies in particular that the concepts $\mathbf{c}_{j}$ are identifiable up to permutation, shift, and scale, and that their representations are identifiable up to a linear transformation. As a result, the concept marginal $p(\mathbf{c})$ is also identifiable (again up to permutation, shift, and scale as well). The shift ambiguity can of course be resolved by assuming the concepts are centered (zero-mean), and the scale ambiguity can be resolved by assuming concepts are normalized (e.g., unit norm).
@@
Assumption 1 (Atomic representations)
There exists a set of atomic representations $\mathcal{A}=\{\mathbf{a}_{1},\ldots,\mathbf{a}_{n}\}$ of linearly independent vectors $\mathbf{a}_{j}\in\mathbb{R}^{d_{e}}$ such that the rows of each representation $C_{j}$ are in $\mathcal{A}$. We denote the indices of $\mathcal{A}$ that appear as rows of $C_{j}$ by $S^{j}$, i.e., $S^{j}=\{i\mathrel{\mathop{\ordinarycolon}}\mathbf{a}_{i}\text{ is a row in $C_{j}$}\}$. We assume that $\cup_{j}S^{j}=\mathcal{A}$, i.e., every atom in $\mathcal{A}$ appears in some concept representation $C_{j}$.
@@
[…]
@@
Assumption 2 (Concept diversity)
@@
B.2 Formal statement of Theorem 4.1
@@
Theorem B.1.
Under Assumptions 1-2 and given single-node interventions on each concept $\mathbf{c}_{j}$, we can identify the representations $C_{j}$ and the latent concept distribution $p(\mathbf{c})$, in the sense of (6-7).
In the context of learning disentangled representations, identifying the concept marginal $p(\mathbf{c})$ is particularly relevant, since this implies we also learn certain (in)dependence relationships between the concepts. In other words, if the “true” concept representations are disentangled (i.e., independent), then we identify disentangled concepts.
@@
  • While our proof ultimately relies on techniques from Rajendran et al. (2024), Theorem B.1 does not trivially follow from their results, which do not cover interventions over the concept marginal $p(\mathbf{c})$. Extending their results to the case of concept interventions, including making sure these interventions are well-defined causal objects, is the main technical difficulty in our proof. This is surprisingly tricky since we do not assume a full structural causal model over $(\mathbf{x},\mathbf{c},\mathbf{e})$: Interventions are only well-defined over $\mathbf{c}$, which makes deducing the downstream effects on the observed $\mathbf{x}$ somewhat subtle.
  • Various results in CRL (e.g. Buchholz et al., 2023; von Kügelgen et al., 2023; Varici et al., 2024, see Section A for more references) prove identifiability results under interventions, however, these results require interventions directly on the embeddings $\mathbf{e}$, as opposed to the concepts $\mathbf{c}$. Theorem B.1, by contrast, requires only $d_{c}\ll d_{e}$ total interventions and in doing so directly addresses the technical challenges associated with intervening directly on concepts as opposed to embeddings.
@@ B.3.1 Concept interventions
Lemma B.3.
For any distribution $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, the interventional distribution $P^{j}$ defined by (11) is also an element of $\mathcal{M}_{\mathbf{c}}(\mathsf{G})$, i.e., $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$.
@@
Proof B.5.
Lemma B.3 implies that $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Thus Lemma B.2 implies the existence of joint distributions $\widetilde{P},\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ such that
[…]
(1) follows from the definition of $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ for $P^{j}\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$ from Lemma B.2 above; i.e., (12) above.
@@
[…]
@@
(3) follows similarly to (2). Namely, since the intervention only perturbs the conditional factors specified above, we can always choose $f^{j}=f$.
@@
Proposition B.6.
Let $I_{j}\mathrel{\mathop{\ordinarycolon}}=\{j\}$ be the collection of single-node intervention targets, and consider the interventional distributions $P^{1},\ldots,P^{d_{c}}$ for $P\in\mathcal{M}_{\mathbf{c}}(\mathsf{G})$. Then there exists $\widetilde{P},\widetilde{P}^{1},\ldots,\widetilde{P}^{d_{c}}\in\mathcal{M}(\mathsf{G})$ such that
@@
Proof B.7.
Fix a choice of $\widetilde{P}\in\mathcal{M}(\mathsf{G})$ for $P$, which specifies a function $f$. For the chosen $\widetilde{P}$, we construct $\widetilde{P},\widetilde{P}^{1},\ldots,\widetilde{P}^{d_{c}}$ according to Proposition B.4. The result follows by applying Proposition B.4 for every $j$.
Proposition B.6 will allow us (below) to define more formally the environments that are used to identify concepts. Intuitively, each environment corresponds to single-node interventions on a single concept $\mathbf{c}_{j}$, however, we can only observe the effect this intervention has on the observed $\mathbf{x}$. The purpose of Propositions B.4-B.6 are to trace the dependence between $\mathbf{x}$ and $\mathbf{c}$—via the embeddings $\mathbf{e}$—according to the model (A1)-(A3).
Formally, let $I_{j}=\{j\}$ with $\eta_{j}^{\prime}\sim\mathcal{N}(0,\omega_{j}^{2})$ and $\mathbf{c}_{j}=\boldsymbol{\mu}_{j}+\eta_{j}^{\prime}$ according to (11), i.e., each environment corresponds to a single-node intervention on the $j$th concept $\mathbf{c}_{j}$. By Proposition B.6, there exist post-intervention joint distributions $\widetilde{P}^{j}\in\mathcal{M}(\mathsf{G})$ over $(\mathbf{x},\mathbf{c},\mathbf{e})$ such that
@@
and $f^{j}=f$ for all $j$. Then the $j$th environment corresponds to sampling $\mathbf{x}=f(\mathbf{e})$ where $\mathbf{e}\sim p_{j}(\mathbf{e})\mathrel{\mathop{\ordinarycolon}}=\widetilde{p}^{j}(\mathbf{e}\,|\,\mathbf{c}_{j}=\boldsymbol{\mu}_{j})$.
B.3.2 Reduction to concept identification
We begin by computing the log-likelihood ratio between the observational and interventional environments. The idea is that the concept representations are revealed through fluctuations in this likelihood ratio. Using Proposition B.6, we have
@@ B.3.2 Reduction to concept identification
$h_{j}(\mathbf{e})\mathrel{\mathop{\ordinarycolon}}=\log p_{0}(\mathbf{e})-\log p_{j}(\mathbf{e})$ $\propto\sum_{i=1}^{n}\frac{1}{2}M_{ij}\mathbf{e}_{i}^{2}-B_{ij}\mathbf{e}_{i}+O(1).$ (16)
These functions are convex, which allows us to identify the $|S_{T}|$ where $S_{T}=\cup_{j\in T}S_{j}$ by minimizing the convex function $\sum_{j\in T}h_{j}(\mathbf{e})$. Then a similar induction argument identifies the matrices $M$ and $B$. Now, given two distinct representations of $p(\mathbf{x})$ via nonlinear mixing functions $f$ and $\widetilde{f}$, define $\varphi=\widetilde{f}^{-1}\circ f$. Let $\mathbf{h}=(h_{1},\ldots,h_{n})$ and write $\mathbf{e}^{c}=(\mathbf{e}_{1},\ldots,\mathbf{e}_{n})$ and $\mathbf{e}=(\mathbf{e}^{c},\mathbf{e}^{\perp})\in\mathbb{R}^{n\times(d_{c}-n)}$. Let $\iota^{\perp}(\mathbf{e}^{c})=(\mathbf{e}^{c},0)$, $\pi^{c}(\mathbf{e})=\mathbf{e}^{c}$, and $\varphi^{\perp}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\to\mathbb{R}^{n}$ by $\varphi^{\perp}(\mathbf{e}^{c})_{i}=\varphi(\mathbf{e},0)_{i}$. Note that via the change of variables formula, we can write $\mathbf{h}(\mathbf{e})=\mathbf{H}(\mathbf{x})$ for some function $\mathbf{H}$. Then
@@
which is the crucial relation used in the proof of Theorem 1 of Rajendran et al. (2024). From here, identifiability follows from a straightforward but lengthy linear algebraic argument based on the first-order conditions for minimizing $\mathbf{h}$ and the identifiability of $M,B$ proved above. The proof of Theorem B.1 follows.
Appendix C quad: A semi-synthetic benchmark for compositional generation
@@
These properties make quad especially useful for both quantitative and qualitative evaluation; indeed, it was created precisely for this reason. Example images are shown in Figures 3-4.
To describe the experimental setting used in the quad (independent) experiments, let $\mathbf{c}_{j}$ indicate the $j$th concept as listed above. We used seven different contexts, including an observational context and one context each for single-concept interventions on $(\mathbf{c}_{1},\ldots,\mathbf{c}_{6})$, i.e., quad1, quad2, quad3, quad4, size, and orientation. We use a sample size of 50 000 images per context. In each context, $\mathbf{c}_{7}$ (object) and $\mathbf{c}_{8}$ (shape) are sampled uniformly at random. The remaining six concepts were sampled as follows:
@@
Similarly, we can compose multiple interventions together, as in Figure 4. This facilitates evaluation of the concept composition capabilities of models trained on single interventions, providing a ground truth distribution to compare against generated OOD samples like in Figure LABEL:fig:quad-ood-full.
Figure 3: Example images obtained from the quad dataset. Each subfigure corresponds to a different context, indicated by the title (obs = observational; the others indicate a single-node intervention). For example, in quad1, the first (top left) quadrant has been manipulated from orange-green hues to blue-red hues.
@@
For comparison, see Figure E.1.2 for examples of generated multi-concept compositions (OOD) from a simple 3-layer convolutional model. The training data did not contain any multi-concept compositions.
Figure 4: Example images obtained from the quad dataset. Each subfigure corresponds to a different double-concept intervention context, indicated by the title. For example, in quad2_quad3, the second (top-right) and third (bottom-left) quadrants have been manipulated from orange-green hues to blue-red hues.
For quad dependent, a random SEM is generated using the sempler package (Gamella et al., 2022), which also allows for data generation under shift interventions.
Appendix D Experimental details
In this appendix we provide additional details on our experimental protocol and set-up. We have used Snakemake (Mölder et al., 2025) for reproducibility.
D.1 Evaluation metrics